### v2

@@ 4.1 Decide and Propagate in Domain
We refer to the set of variables that need to be considered during solving as the domain. For solving $Q_{relind}$, it is sufficient to use $\mathcal{V}(F_{i})\cup\mathcal{V}(c)\cup COI(c^{\prime})$ as the domain, instead of all variables. This significantly reduces the number of variables that the SAT solver needs to decide and assign during BCP. We need to efficiently compute the domain before each solving process. For $\mathcal{V}(F_{i})$, we maintain a data structure within the SAT solver that tracks and updates it whenever a new lemma is added. For $\mathcal{V}(c)$, we simply extract all the variables that appear in $c$. Finally, $COI(c^{\prime})$ is computed by recursively traversing the fanins of $c^{\prime}$.
At the beginning of the solving process, we push all variables from the computed domain into the VSIDS decision-required set, ensuring that all domain variables will be decided. During BCP, if a clause includes a variable not in the current domain, we skip that clause and move on to the next one. This approach ensures that variables outside the domain are not assigned during BCP.
@@ 6.2 Experimental Results
Configuration #Solved $\Delta$Solved Avg. ST Ratio PAR-2
rIC3-GipSAT 392 0 x1.00 2843.50
@@
rIC3-CryptoMinisat 364 -28 x4.83 3216.31
rIC3-CaDiCaL 366 -26 x5.99 3189.09
ABC 357 - - 3258.97
nuXmv 353 - - 3289.72
@@
Figure 5: The comparison of solving time(s) between GipSAT and other solvers in rIC3. Points below the gray solid line indicate that GipSAT performs better. The purple dashed line represents the geometric mean of the speedup.
@@ 6.3 Analysis
To better demonstrate the effectiveness of GipSAT, we measured the average solving time of queries using GipSAT and Minisat in rIC3, as described in Fig. 6. It is evident that, in the majority of cases, GipSAT demonstrates a significant reduction in solving time for queries when compared to Minisat. We also evaluated rIC3-GipSAT under different conditions: when it makes decisions and propagates in all variables instead of the domain (rIC3-GipSAT-av), when it utilizes a binary heap instead of buckets in VSIDS (rIC3-GipSAT-bh), and when it creates a new activation variable before each solving and resets the solver every 1000 activation variables (rIC3-GipSAT-rs). The results, which are presented in Table 3, demonstrate the effectiveness of deciding in the domain, using buckets in VSIDS, and eliminating resetting respectively.
@@
Table 3: The total number of cases of different conditions. Using rIC3-GipSAT as the baseline, the additional solved cases and the geometric mean of the solving time ratio compared to the baseline.
Configuration #Solved $\Delta$Solved Avg. ST Ratio
rIC3-GipSAT 392 0 x1.00
@@
rIC3-GipSAT-bh 386 -6 x1.29
rIC3-GipSAT-rs 390 -2 x1.14
@@ 7 Related Work
Since the introduction of IC3 [12], numerous efforts have been dedicated to enhancing its performance. Most of these studies focus on the algorithm layer. For instance, [25] attempts to block counterexamples to generalization (CTG) to improve generalization. [36] proposes a variant of IC3 that integrates interpolation, while [31] introduces under-approximation to expedite bug detection. The algorithm in [38] drops literals that do not appear in any subsumed lemmas from the previous frame, thereby increasing the likelihood of propagating to the next frame. In [34], the authors aim to predict the outcome before generalization, which could potentially reduce overhead if successful. Additionally, [24] aggressively pushes lemmas to the top by adding may-proof-obligation.
Only a few studies have focused on the SAT solver layer. [14] analyzed the results of different SAT solver management strategies at the algorithm layer, including SAT solver allocation, loading, and clean-up. [21] proposed a method that eliminates the need for activation literals in temporary clauses. Instead, it iteratively assumes a literal in the temporary clause before solving, thereby avoiding the need to reset the SAT solver. Similarly, GipSAT also eliminates the need for resetting the SAT solver, but it differs significantly in its approach. GipSAT uses the same activation literal in every query by promptly removing both the temporary clause and its derived learned clauses after each solving.

### v3

@@ 4.1 Decide and Propagate in Domain
We refer to the set of variables that need to be considered during solving as the domain. For solving $Q_{relind}$, it is sufficient to use $\mathcal{V}(F_{i})\cup\mathcal{V}(c)\cup COI(c^{\prime})$ as the domain, instead of all variables. This significantly reduces the number of variables that the SAT solver needs to decide and assign during BCP.
We use a hash table to record whether each variable belongs to the domain, allowing constant-time queries for domain information. Before each solving process, we need to compute the domain relevant to that instance. For $\mathcal{V}(F_{i})$, the variables are permanently stored in the hash table, and new domain variables are added whenever a new lemma is introduced. For $\mathcal{V}(c)$, we simply extract all variables that appear in clause $c$. Finally, $COI(c^{\prime})$ is computed by recursively traversing the fanins of $c^{\prime}$. After computing $\mathcal{V}(c)$ and $COI(c^{\prime})$, we also temporarily record them in the domain hash table, but they are removed after each solving process.
After computing the domain for the current solving instance, we push all domain variables into the VSIDS decision-required set to ensure they are considered for decisions. During BCP, if a clause contains a variable that is not in the current domain, it is considered irrelevant to the current solving process. Since out-of-domain variables are not eligible for assignment, any clause involving them cannot contribute to further propagation. When BCP attempts to find a new unassigned variable to serve as the watch literal, if that variable lies outside the domain, we simply ignore the clause and proceed to the next one, without performing any variable assignments based on that clause. This approach guarantees that out-of-domain variables are never assigned during BCP. Furthermore, by restricting assignments to domain variables only, we eliminate propagation involving irrelevant variables, thereby reducing BCP overhead.
@@ 5 Implementation
  • inductive_core(): The inductive core is the UNSAT core produced by the previous relind($c$), and an extra condition ensures that it does not overlap with the initial state $I$.
@@ 6.2 Experimental Results
Configuration #Solved $\Delta$Solved #Safe #Unsafe Avg. ST Ratio PAR-2
rIC3-GipSAT 392 0 326 66 x1.00 2843.50
@@
rIC3-CryptoMinisat 364 -28 310 54 x4.83 3216.31
rIC3-CaDiCaL 366 -26 311 55 x5.99 3189.09
ABC 357 - 305 52 - 3258.97
nuXmv 353 - 301 52 - 3289.72
@@
Figure 5: The comparison of solving time(s) between GipSAT and other solvers in rIC3. Points above the gray solid line indicate that GipSAT performs better. The purple dashed line represents the geometric mean of the speedup.
@@ 6.3 Analysis
To better demonstrate the effectiveness of GipSAT, we measured the average solving time of queries using GipSAT and Minisat in rIC3, as described in Fig. 6. It is evident that, in the majority of cases, GipSAT demonstrates a significant reduction in solving time for queries when compared to Minisat. We also compared the proof-obligation length, number of frames, and invariant size between GipSAT and Minisat, as shown in Fig. 7. For cases solved by both GipSAT and Minisat, these values are largely the same from a statistical perspective, indicating that our approach primarily accelerates SAT solving without significantly affecting the algorithm itself. For cases where GipSAT and Minisat do not both solve the problem, the frame depth of GipSAT is slightly larger. This is because GipSAT’s faster solving speed allows the algorithm to reach deeper frames more quickly within the given time limit.
@@
Figure 7: The comparison of proof-obligation length, number of frames, and invariant size (safe instances) between GipSAT and Minisat in rIC3.
To assess the impact of each optimization, we evaluated rIC3-GipSAT under different optimization combinations:
  • The dm flag determines whether domain optimization is enabled.
  • The bkt flag indicates whether buckets are used to maintain VSIDS; if not, a binary heap is used instead.
  • The wr flag represents whether the optimization described in Section 4.3 is applied; otherwise, a new activation variable is created before each solving, and the solver is reset every 1000 activation variables.
The results are presented in Table 3, where we also include the results of rIC3-Minisat for comparison. In the table, rIC3-GipSAT-noopt refers to the version with all optimizations disabled. As shown, without any optimizations, rIC3-GipSAT-noopt performs similarly to rIC3-Minisat as GipSAT’s CDCL framework is inspired by Minisat. Additionally, it can be observed that enabling optimizations leads to performance improvements, demonstrating the effectiveness of deciding in the domain, using buckets in VSIDS, and eliminating resetting, respectively. For further comparison, we also implemented the method from [22] on GipSAT (rIC3-GipSAT-fmcad21), which performs slightly better than ours. This method sequentially treats each literal in the constraint as an assumption, resulting in support for only one temporary clause, whereas our method allows for multiple temporary clauses.
7 Discussion
The IC3 algorithm involves a large number of relatively simple SAT problems, so improving the efficiency of each SAT call is essential for enhancing overall performance. The first optimization significantly boosts performance by reducing the number of decision variables, which leads to substantial improvements in many cases. However, when domain variables make up a large portion of the total variables, the effectiveness of this optimization becomes limited. The second optimization replaces the binary heap with a bucket-based approach to reduce the overhead of the VSIDS heuristic, while the third reuses activation literals to avoid unnecessary resets. Nonetheless, compared to the first optimization, the performance gains from the second and third optimizations are relatively modest. Overall, GipSAT significantly improves the SAT solving efficiency of the IC3 algorithm, enabling more cases to be solved within a limited time. However, it does not alter the IC3 algorithm itself, so cases that are unsolvable due to theoretical limitations remain unsolved.
Table 3: The total number of cases of different conditions. Using rIC3-GipSAT-noopt as the baseline, the additional solved cases and the geometric mean of the solving time improvement compared to the baseline are presented.
Configuration #Solved $\Delta$Solved Avg. ST Imprv
rIC3-GipSAT-noopt 370 0 x1.00
@@ 7 Discussion
rIC3-GipSAT-dm 384 +14 x2.62
rIC3-GipSAT-bkt 374 +4 x1.24
rIC3-GipSAT-wr 373 +3 x1.13
rIC3-GipSAT-fmcad21 373 +3 x1.21
rIC3-GipSAT-bkt-wr 379 +9 x1.42
rIC3-GipSAT-dm-wr 386 +16 x2.84
rIC3-GipSAT-dm-bkt 390 +20 x3.17
rIC3-GipSAT-dm-bkt-wr 392 +22 x3.52
@@ 8 Related Work
Since the introduction of IC3 [12], numerous efforts have been dedicated to enhancing its performance. Most of these studies focus on the algorithm layer. For instance, [27] attempts to block counterexamples to generalization (CTG) to improve generalization. [38] proposes a variant of IC3 that integrates interpolation, while [33] introduces under-approximation to expedite bug detection. The algorithm in [40] drops literals that do not appear in any subsumed lemmas from the previous frame, thereby increasing the likelihood of propagating to the next frame. In [36], the authors aim to predict the outcome before generalization, which could potentially reduce overhead if successful. In [18], the authors leverage the internal signals of the circuit to represent invariants more concisely. [26] presents a flexible algorithmic framework that integrates IC3 with k-induction. Additionally, [25] aggressively pushes lemmas to the top by adding may-proof-obligation.
Only a few studies have focused on the SAT solver layer. [14] analyzed the results of different SAT solver management strategies at the algorithm layer, including SAT solver allocation, loading, and clean-up. [22] proposed a method that eliminates the need for activation literals in temporary clauses. Instead, it iteratively assumes a literal in the temporary clause before solving, thereby avoiding the need to reset the SAT solver, which means it can only support one temporary clause. Similarly, GipSAT also eliminates the need to reset the SAT solver, but it differs significantly in its approach. GipSAT uses the same activation literal in every query by promptly removing both the temporary clause and its derived learned clauses after each solving, allowing multiple temporary clauses. In Minisat, the release_var interface can avoid restarts by passing the activation variables. The principle behind it is to first add the activation variable clause and then periodically clean up satisfied clauses (since the occur list is not maintained for performance reasons, it can only be lazily cleaned up). Once all clauses containing released variables are cleared, the activation variables can be reused. Compared to this approach, our method not only avoids restarts but also promptly removes temporary clauses. In SMT, temporary constraints can be supported through push and pop, but more complex data structures are needed to maintain them.
@@ 9 Conclusion and Future Work
Acknowledgement
This work was supported by the Beijing Municipal Natural Science Foundation (Grant No. 4252024), the Foundation of Laboratory for Advanced Computing and Intelligence Engineering (Grant No. 2023-LYJJ-01-013), the Basic Research Projects from the Institute of Software, Chinese Academy of Sciences (Grant No. ISCAS-JCZD-202307) and the National Natural Science Foundation of China (Grant No. 62372438).
Disclosure of Interests
The authors have no competing interests to declare that are relevant to the content of this article.